### v2

# Contrastive Self-Supervised Learning As Neural Manifold Packing

Abstract
Contrastive self-supervised learning based on point-wise comparisons has been widely studied for vision tasks. In the visual cortex of the brain, neuronal responses to distinct stimulus classes are organized into geometric structures known as neural manifolds. Accurate classification of stimuli can be achieved by effectively separating these manifolds, akin to solving a packing problem. We introduce Contrastive Learning As Manifold Packing (CLAMP), a self-supervised framework that recasts representation learning as a manifold packing problem. CLAMP introduces a loss function inspired by the potential energy of short-range repulsive particle systems, such as those encountered in the physics of simple liquids and jammed packings. In this framework, each class consists of sub‑manifolds embedding multiple augmented views of a single image. The sizes and positions of the sub-manifolds are dynamically optimized by following the gradient of a packing loss. This approach yields interpretable dynamics in the embedding space that parallel jamming physics, and introduces geometrically meaningful hyperparameters within the loss function. Under the standard linear evaluation protocol, which freezes the backbone and trains only a linear classifier, CLAMP achieves competitive performance with state‑of‑the‑art self‑supervised models. Furthermore, our analysis reveals that neural manifolds corresponding to different categories emerge naturally and are effectively separated in the learned representation space, highlighting the potential of CLAMP to bridge insights from physics, neural science, and machine learning.
1 Introduction
Learning image representations that are both robust and broadly transferable remains a major challenge in computer vision, with critical implications for the efficiency and accuracy of training across downstream tasks. State-of-the-art models address this problem by employing contrastive self-supervised learning (SSL), wherein embeddings are treated as vectors and optimized using pairwise losses to draw together positive samples, augmented views of the same image, and push apart negative samples, views of different images [1, 2, 3, 4, 5, 6]. While SSL has advanced significantly, and in some cases, surpassed supervised methods [7, 8], its geometric foundations remain largely underexplored. In visual cortex, neural activity often resides within geometric structures known as neural manifolds — which span the embedding space [9] and may exhibit low intrinsic dimensionality [10] that capture much of the observed variability in neural activity. Neural manifolds have been characterized in a number of neural systems supporting a variety of cognitive processes (in addition to vision) including neural activity coordination, motor control, and information coding [11, 12, 13, 14]. In the context of SSL, neural representations of similar and dissimilar samples are frequently conceptualized as distinct sub-manifolds in the representation space, where a sub-manifold corresponds to the encoding of a single data sample and its transformations, and a manifold to an entire class. In classification, the objective is to separate these class-level manifolds so that each maps cleanly to a distinct category under a readout function. When manifolds corresponding to different classes, for example, manifolds representing “cats” and “dogs” overlap in neural state space, the resulting ambiguity can lead to misclassification. This viewpoint recasts classification as the problem of efficiently arranging manifolds in representation space to minimize overlap and enhance separability [15].
Our main contributions are as follows:
  • We propose a loss function inspired by the interaction energy of short-range repulsive particle systems such as simple liquids and jammed packings, endowing each hyperparameter with a clear physical interpretation.
  • We demonstrate that CLAMP achieves state-of-the-art image classification accuracy, matching state-of-the-art SSL methods and even setting a new SOTA on ImageNet-100, under the standard linear evaluation protocol (a frozen backbone plus a linear readout), all while being remarkably simple and interpretable thanks to its single projection head and one-term, physics-grounded loss comprising a single term.
  • We show that the dynamics of SSL pretraining under CLAMP resemble those of the corresponding interacting particle systems, revealing a connection between SSL and non-equilibrium physics.
  • We compare CLAMP’s internal representations to mammalian visual cortex recordings and find that CLAMP offers a compelling account of how complex visual features may self-organize in the cortex without explicit supervision.
2 Related work
2.1 Interacting particle models in physics
A particularly insightful analogy for CLAMP comes from random organization models, originally introduced to describe the dynamics of driven colloidal suspensions [16, 17, 18]. In these models, overlapping particles are randomly “kicked” while isolated particles remain static. As the volume fraction is increased, the system undergoes a phase transition from an absorbing state (where no particles overlap) to an active steady state (where overlaps cannot be resolved under the prescribed dynamics) in two and three dimensions [19]. When kicks between overlapping particle pairs are reciprocal, meaning that each particle is displaced by an equal but opposite amount, the critical and active states are characterized by anomalously suppressed density fluctuations at large scales, a property known as “hyperuniformity”[20] or “blue noise”[21, 22]. This absorbing-to-active transition coincides with random close packing (RCP) when the noisy dynamics are reciprocal [17]. Recent work further establishes an equivalence between this process and a stochastic update scheme that minimizes a repulsive energy under multiplicative noise [15]. We note that even beyond the hyperuniformity observed for dense active states of random organization systems, for a high-dimensional system of short-range, repulsive particles, such as those interacting via an exponential potential ($e^{-cr},c>0$), confined to a closed surface, a uniform spatial distribution minimizes the system’s energy in the large density limit [23, 24, 25]. Drawing on these insights, CLAMP employs a short-range repulsive potential over augmentation sub-manifolds as its self-supervised learning loss, favoring the separation of embeddings.
2.2 Feature-level contrastive learning
Following recent SOTA self-supervised learning models, we use image augmentations and contrastive losses. Most self-supervised learning frameworks treat features as points or vectors in the embedding space, contrasting positive samples—embeddings of augmented views of the same image—with negative samples. They typically use a pairwise contrastive loss such as InfoNCE [1, 2, 3, 4, 5, 6], or a weighted-sample InfoNCE loss where negative samples are drawn from tailored Boltzmann-like distributions to sharpen sample-wise contrasts [26]. Here, CLAMP treats each image as its own class; however, instead of applying the contrastive loss directly to individual feature vectors, it operates on the entire augmentation sub-manifolds surrounding each input embedding.
2.3 Clustering methods for self-supervised learning
Clustering-based self-supervised learning methods assign images to clusters based on similarity criteria, rather than treating each image as a distinct class. For instance, The DeepCluster [27] applies the k-means algorithm to assign pseudo-labels to embeddings, which serve as training targets in subsequent epochs. Collapse-resistant non-contrastive methods prevent feature collapse by using a fixed dictionary and enforcing augmentation invariance without negative samples [28]. However, the DeepCluster approach requires k-means classification on the entire dataset, making it unsuitable for online, mini-batch, learning. SwAV [29] overcomes this limitation by using soft assignments and jointly learning cluster prototypes and embeddings, enabling flexible online updates. In contrast to clustering methods, our approach groups each image with its augmentations to learn the geometry of the augmentation sub-manifold without clustering across different samples.
2.4 Alignment-Uniformity and maximum manifold capacity methods
The Alignment-Uniformity method [30] constructs the loss function by incorporating two components: an alignment term and a uniformity term. The alignment term maximizes the cosine similarity of the embedding vectors between positive pairs, thereby encouraging consistency. The uniformity term acts as a repulsive pairwise potential over all pairs of embeddings, driving the representations to be uniformly distributed on the hyper-sphere and mitigating representational collapse [31].
Maximum manifold capacity theory measures the linear binary classification capacity of manifolds with random binary labels. For $P$ $d$-dimensional manifolds embedded in $D$ dimensions, $\alpha_{c}$ measures the maximum capacity such that for all $P/D<\alpha_{c}$, these manifolds can be linearly classified into two categories with probability 1 in the limit where $P\rightarrow\infty$, $D\rightarrow\infty$, $d/D\to 0$, and $P/D$ remains finite [32, 33]. Maximum manifold capacity representation (MMCR) maximizes this capacity on augmentation sub-manifolds in the framework of multi-view self-supervised learning and achieves high accuracy classification performance [34].
Unlike MMCR, which expressly maximizes the representation’s binary linear classification capacity, CLAMP optimizes the embeddings for an $n$-ary nonlinear classification problem, making the two approaches fundamentally different. Moreover, CLAMP’s loss function is based on short-range interactions: the contribution from each augmentation sub-manifold depends only on its neighboring sub-manifolds, whereas MMCR requires global information from all sub-manifolds. Finally, although MMCR has comparable computational complexity to CLAMP, MMCR is less efficient to train, because it requires one SVD for each learning step, compared to CLAMP, and its loss function and optimization dynamics lack the clear, physics-based interpretability that CLAMP provides.
Compared to the Alignment-Uniformity method, our approach differs in three key aspects: (1) The Alignment-Uniformity loss is defined over feature vectors, whereas our loss operates directly on manifolds; (2) Our method employs a single loss term that simultaneously enforces alignment and uniformity, avoiding the need to manually balance separate terms as required in Alignment-Uniformity; (3) Our pairwise repulsion is short-ranged, involving only neighboring embeddings, which reduces computational complexity compared to the global (long-ranged) repulsion used in Alignment-Uniformity.
To summarize, both the MMCR and Alignment-Uniformity methods enhance mutual information between positive pairs by pulling together augmentations of the same image (positive samples) and pushing apart those of different images (negative samples) [35, 30]. In CLAMP, these properties emerge naturally from the short-range repulsive dynamics described in Sec. 2.1, without relying on global information or separate alignment and uniformity terms that need to be balanced, making CLAMP a particularly elegant, efficient, and interpretable solution.
3 Method
3.1 Description
Inspired by physical packing problems and recent advances in self-supervised learning, CLAMP learns structured representations by minimizing an energy-based loss that encourages optimal packing of neural manifolds in embedding space. Our framework is designed specifically for vision tasks that use images as input. As shown in Fig. 1, given a batch of images, denoted $x_{1},x_{2},…,x_{b}$, where $b$ represents the batch size, we adopt the standard contrastive learning setup, in which positive samples are generated through image augmentations. For each input image $x_{i}$, we apply random transformations to obtain $m$ augmented views, denoted as $x_{i}^{1},x_{i}^{2},…,x_{i}^{m}$.
To establish precise terminology, we define an augmentation sub-manifold as the set of embeddings produced from a finite number of augmentations of a single image. In the limit of infinite augmentations, this becomes a sub-manifold. The union of sub-manifolds corresponding to all images in a given class defines a class manifold, or simply a neural manifold.
Figure 1: CLAMP architecture. The CLAMP framework processes a batch of $b$ input images by applying augmentations to generate $m$ views for each image. These augmented views are then encoded and projected into a shared embedding space. Within this space, the augmented embeddings corresponding to each input form a distinct sub-manifold, resulting in $b$ such sub-manifolds. Then, a pairwise packing loss is applied to minimize overlap between these sub-manifolds. The gradient of the loss is subsequently backpropagated to optimize the model.
[…]
For each image $x_{i}$, its $m$ augmentations form a cluster of projected vectors that we refer to as an augmentation sub-manifold, $\{\tilde{z}_{i}^{1},\tilde{z}_{i}^{2},...\tilde{z}_{i}^{m}\}$ centered at $Z_{i}=\frac{1}{m}\sum_{k}\tilde{z}_{i}^{k}$. We approximate the shape of each augmentation sub-manifold with a high-dimensional ellipsoid, as described next in Sec. 3.1.1. This procedure yields $b$ ellipsoidal sub-manifolds, one per image. We treat sub-manifolds from different images as repelling entities and define the loss as $\mathcal{L}=\log(\mathcal{L}_{overlap})$, where $\mathcal{L}_{overlap}=\sum_{i,j}E(Z_{i},Z_{j})$ penalizes overlap between sub-manifolds. $E(Z_{i},Z_{j})$ is an energy-based pairwise repulsive potential that pushes ellipsoids $i$ and $j$ away from one another if they overlap,
$\begin{split}&\mathcal{L}_{overlap}=\begin{cases}\sum\limits_{i\neq j}\left(1-\frac{\lVert Z_{i}-Z_{j}\rVert_{2}}{r_{i}+r_{j}}\right)^{2}\text{, if }\lVert Z_{i}-Z_{j}\rVert_{2}<r_{i}+r_{j}\\ 0\text{,\qquad otherwise }\end{cases},\text{where }r_{i}=r_{s}\sqrt{\frac{\mathrm{Tr}(\Lambda_{i})}{m}}.\end{split}$ (1)
Note that even though the loss function acts locally on neighboring sub-manifolds in the embedding space, each weight update affects all embeddings. Since the logarithm is a monotonic function, applying it to $\mathcal{L}_{\text{overlap}}$ preserves the locations of its minima while compressing large loss values, stabilizing the training process.
Following the convention established in [2], we refer to the output space of the encoder $f_{\theta}(\cdot)$ as the representation space, while the output space of the projection head $g_{\phi}(f_{\theta}(\cdot))$ is designated as the embedding space.
3.1.1 Approximate augmentation sub-manifolds as ellipsoids
We approximate the distribution of embedding vectors $p(\tilde{z}_{i})$ generated from random augmentations of input image $x_{i}$ as a multivariate Gaussian with mean $\tilde{Z}_{i}$ and covariance $\Lambda_{i}$, $\tilde{z}_{i}\sim\mathcal{N}(Z_{i},\Lambda_{i})$. We assume that $\Lambda_{i}$ is full-rank and invertible (see Appendix K for a treatment of the general case). To represent the augmentation sub-manifold in embedding space, we define the corresponding enclosing ellipsoid with Mahalanobis distance $r_{s}$,
$(\tilde{z}-Z_{i})\Lambda_{i}^{-1}(\tilde{z}-Z_{i})^{T}=r_{s}^{2}$ (2)
where $r_{s}$ is a scaling hyperparameter that controls the effective size of the manifold. A large $r_{s}$ corresponds to larger ellipsoids, and an appropriate choice of $r_{s}$ ensures that the ellipsoid encloses the majority of the augmentation sub-manifold (Fig. 2(a)). The lengths of the ellipsoid’s semi-axes are given by the square roots of the nonzero eigenvalues of $\Lambda_{i}$ scaled by $r_{s}$: $r_{s}\sqrt{\lambda_{i}^{(1)}},r_{s}\sqrt{\lambda_{i}^{(2)}},$... where $\lambda_{i}^{(k)}$ is the k-th eigenvalue of $\Lambda_{i}$. To estimate the effective size of the ellipsoid, we compute the square root of the expected squared semi-axial length, $\sqrt{\mathbb{E}[(r_{s}\sqrt{\lambda})^{2}]}$, yielding the sub-manifold radius, $r_{i}=r_{s}\sqrt{\frac{\sum_{j}\lambda_{i}^{(j)}}{\text{rank}(\Lambda_{i})}}$. The radius $r_{i}$ provides an upper bound on the volume of the ellipsoid $i$ ($V_{i}$), $r_{i}\geq cV_{i}^{1/\min(b,D)}$ (see Appendix K for details). In practice, the number of augmentations $m\sim O(10)$ is much smaller than the embedding dimension $D\sim O(100)$, making the empirical covariance $\Lambda_{i}$ lower rank or singular. In such cases, the expression for $r_{i}$ remains valid with $\mathrm{rank}(\Lambda_{i})\approx m$ serving as a proxy (see Appendix K for discussion).
3.1.2 Pairwise potential enhances similarity between positive samples and prevents embeddings from collapsing
CLAMP models sub-manifolds as repulsive particles interacting via a short-range potential $E(Z_{i},Z_{j})$, where $Z_{i},Z_{j}$ are the centroids of sub-manifolds $i$ and $j$. The potential attains its maximum at $Z_{i}-Z_{j}||_{2}=0$ corresponding to complete overlap between two sub-manifolds and drops to zero when $Z_{i}-Z_{j}||_{2}>r_{i}+r_{j}$, ensuring that non-overlapping sub-manifolds do not interact. Eq. 1 implies that increasing the inter-center distance between sub-manifolds and reducing their sizes leads to a lower loss.
Similarity. Similarity requires that augmentations of the same image (positive samples) be mapped to nearby points in the feature space, thereby making them robust to irrelevant variations. CLAMP allows the manifold sizes to vary dynamically, unlike fixed-size particle-based models in physics. This flexibility is crucial in self-supervised representation learning, where small augmentation sub-manifold sizes $r_{i}$ are desirable to enhance similarity across positive samples.
Avoiding representational collapse. Representational collapse refers to a degenerate solution in which all network outputs converge to a constant vector, regardless of the input [31]. In CLAMP, the repulsive term in the loss function avoids such collapse by explicitly pushing sub-manifolds of different images apart.
Separability. Separability encourages the negative embeddings to be distant and separable on the unit hypersphere. Analogous repulsive interactions in physical systems are known to produce near-uniform distributions. Consistent with this, we observe that the embeddings progressively become more separable throughout training. To quantify this effect, we track both similarity (distance among positives) and separability (distance among negatives) throughout pretraining, confirming that CLAMP systematically enhances class-level separation.
Figure 2: Sub-manifold and visualization of the embedding space. (a) Schematic for approximating augmentation sub-manifolds as ellipsoids with different scale factor $r_{s}$. (b) We selected 10 images from the MNIST dataset, one for each digit from 0 to 9, and applied Gaussian noise augmentation. These augmented images were then encoded into a 3-dimensional embedding space for visualization. Solid dots represent the embedding points of each augmented view, while the shaded regions denote circumscribed ellipsoids defined by $(\tilde{z}-Z_{i})(\Lambda_{i})^{-1}(\tilde{z}-Z_{i})=r_{s}^{2}$. Left: the initial embeddings. Right: the trained embeddings. For this toy example, we use $r_{s}=3.0$
To visualize CLAMP and provide a proof of concept, we developed a toy model using an 18-layer ResNet as the backbone, $f_{\theta}$, and an MLP, $g_{\theta}$, that maps the resulting representation to three-dimensional embedding space. We loaded one image per class from the MNIST dateset [37] and applied Gaussian noise to generate 60 augmented views per image, and set $r_{s}=3.0$. Fig.2(b) shows the manifold repulsion and size shrinkage dynamics of CLAMP.
3.2 Implementation details
The code for CLAMP is available at https://github.com/guanming-zhang/clamp
3.2.1 Image augmentation
During data loading, each image was transformed into $m$ augmented views using asymmetric augmentation consisting of two augmentation pipelines each sampled from different transformation distributions and applied to one half of the views. We used the following augmentations: random cropping, resizing, random horizontal flipping, random color jittering, random grayscale conversion, random Gaussian blurring (see Appendix B for details).
3.2.2 Architecture
We used a ResNet network, $f_{\theta}(x)$, as the backbone and a MLP, $g_{\phi}(x)$, as the projection head. Following [1] and [2], the projection head MLP was discarded after pretraining.
CIFAR10  For the CIFAR10 dataset [38], we used the ResNet-18 [39] network as the backbone and a two-layer MLP of size [2048,128] as the projection head. Following [4, 2], we removed the max pooling layer and modified the first convolution layer to be kernel_size=1,stride=1,padding=2.
ImageNet-100/1K  For the ImageNet dataset [40], we used the ResNet-50 [39] network as the backbone and a three-layer MLP of size [8192,8192,512] as the projection head.
3.2.3 Optimization
We used the LARS optimizer [41] for training the network. For CIFAR10, we trained the model for 1000 epochs using the warmup-decay learning rate scheduler, and with 10 warmup steps and a cosine decay for rest of the steps. For ImageNet-100, we trained the model for 200 epochs with 10 steps of warmup and 190 steps of cosine decay for the learning rate. For ImageNet-1K, we trained the model for 100 epochs with 10 steps of warmup and 90 steps of cosine decay for the learning rate. Training on 8 A100 GPUs with batch size 1024 for 4 views with distributed data parallelization for 100 epochs took approximately 17 hours. See Appendix D for details.
Linear evaluation Semi-supervised
  Method ImageNet-100 ImageNet-1K 1% 10%
  SimCLR[1] 79.64 66.5 42.6 61.6
SwAV[29] - 72.1 49.8 66.9
Barlow Twins[42] 80.38* 68.7 45.1 61.7
BYOL[2] 80.32* 69.3 49.8 65.0
VICReg[6] 79.4 68.7 44.75 62.16
CorInfoMax[43] 80.48 69.08 44.89 64.36
MoCo-V2[3] 79.28* 67.4 43.4 63.2
SimSiam[4] 81.6 68.1 - -
A&U [30] 74.6 67.69 - -
MMCR (4 views+ME)[34] 82.88 71.5 49.4 66.0
  CLAMP (4 views) 85.12 $\pm.05$ 69.50 $\pm.14$ 47.38 $\pm.56$ 65.10 $\pm.30$
CLAMP (8 views) 85.10 $\pm.15$ 70.04 $\pm.16$ 47.87 $\pm.03$ 65.96 $\pm.04$
Table 1: Linear evaluation accuracy on ImageNet-100 dataset for 200-epoch pretraining: For SimCLR the higher accuracy between [34] and [43] is reported. For Barlow twins, BYOL, and MoCo-V2, the accuracies using ResNet-18 as the backbone (labeled with *) in [43] are reported because either ResNet-50 results are absent or ResNet-18 performs better in [43]. A&U and MMCR results are taken from [30] and [34] respectively. Linear evaluation accuracy on ImageNet-1K dataset for 100-epoch pretraining: For SimCLR, SWAV, Barlow Twins, BYOL, VICReg, CorInfoMax, MoCo-V2 and SimSiam, we reported the highest accuracy for each method among the results in the solo-learn library [44], VISSL library [45], [34] and [43]. We report the 200-epoch pretraining result for A&U, as the 100-epoch result is unavailable [30]. Semi-supervised learning:The semi-supervised learning results for VICReg and CorInfoMax are taken from [43], those for other methods are from [34]. In MMCR with 4 views and a momentum encoder (4 views + ME), the effective 8 views, 4 from the backbone and 4 from the encoder, achieve the same linear accuracy (71.5%) as using 8 views without the momentum encoder [34]
4 Evaluation
4.1 Linear evaluation
Following standard linear evaluation protocols, we froze the pretrained ResNet-50 backbone encoder and trained a linear classifier on top of the representation. Training was conducted for 100 epochs on ImageNet-1K and 200 epochs on ImageNet-100. Classification accuracies are reported on the corresponding validation sets (Table 1). We find that CLAMP achieves competitive performance relative to existing SSL algorithms and even sets a new state of the art on ImageNet-100. Our model performance under linear evaluation depends only weakly on batch size but improves with an increased number of augmented views as shown in Appendix L. See Appendix E for details about linear evaluation.
4.2 Semi-supervised learning
We evaluated the model using a semi-supervised setup with $1\%$ and $10\%$ split (the same as [1]) of the ImageNet-1K training dataset. During semi-supervised learning, both the backbone encoder and the appended linear classifier were updated for 20 epochs. The results are shown in Table 1. Again, we find that CLAMP achieves competitive performance relative to existing SSL algorithms. See Appendix G for details.
4.3 Transfer to object detection tasks
To evaluate CLAMP’s generalizability beyond linear classification, we apply the pretrained model to object detection. Specifically, we report results using a Faster R-CNN architecture with a C4 backbone, pretrained with CLAMP on ImageNet-1K using 8 views and a batch size of 512 for 100 epochs. The model is fine-tuned on the VOC2007+2012 training set and evaluated on the VOC2007 test set (Table. 2). CLAMP delivers stronger detection performance than other self-supervised learning baselines, demonstrating that its learned manifold structure transfers effectively to spatially structured vision tasks. See Appendix H for details.
Methods mAP $\uparrow$ AP50 $\uparrow$ AP75 $\uparrow$
  SimCLR 54.4 81.6 61.0
Barlow Twins 53.1 80.9 57.7
BYOL 55.6 82.3 62.0
MoCo v2 54.7 81.7 60.2
MMCR 54.6 81.9 60.0
CLAMP 55.7 82.3 62.4
Table 2: Object detection on VOC datasets: The benchmarks (mAP, AP50 and AP75) for the baseline methods are from [34]. Models are pretrained on ImageNet-1K for 100 epochs before fine-tuning.
5 Training dynamics reflect geometrical changes in the embedding space
We analyze manifold packing dynamics by tracking the evolution of neighbor count and manifold size during ImageNet-1K training. Two sub-manifolds are considered neighbors if the Euclidean distance between their centroids is smaller than the sum of their radii ($q_{i}-q_{j}||2<r_{i}+r_{j}$). At each epoch, we compute the average number of neighbors (for $m=4$ views) and the average manifold size, $\mathbb{E}_{i\sim\mathrm{data}}[\sqrt{\Lambda_{i}}/m]$, using a 1% validation split. As shown in Fig. 3, both metrics decrease over training, indicating that embeddings evolve from an initial collapsed state toward more structured representations with increased pairwise distances. This mirrors random organization models [17, 15], where local density and spacing show similar temporal behavior, reflecting improved feature discrimination over time.
Figure 3: Training dynamics: (a) Number of neighbours as a function of epochs. (b) Average embedding sub-manifold sizes as the function of epochs. (c) Distances between pair of embeddings for untrained and trained networks.
Figure 4: The properties of sub-manifolds in the embedding space for the pretrained ResNet-18 network are characterized by: (a) Orientation similarity: the squared cosine similarity between the principal orientations of the sub-manifolds. (b) Centroid distances: the Euclidean distances between the centroids of different sub-manifolds. (c) Centroid similarity: the cosine similarity between the centroid points of sub-manifolds.
6 Properties of the representations
Figure 5: t-SNE visualization of the representations. Visualization of the 256-dimensional representation space by t-SNE method. Each color shows the representation corresponding to different category in CIFAR-10 dataset.
To quantify the geometric properties of the learned representations, we analyzed the CLAMP representation obtained for CIFAR-10 using a ResNet-18 backbone architecture. As shown in Appendix F, we achieve $\sim 90\%$ top-1 accuracy (without hyperparameter tuning) which is competitive with alternative SSL frameworks. For each sub-manifold, we computed its centroid and alignment vector (defined as the principal eigenvector of its covariance matrix), grouping them according to their class labels. We compared the geometric properties such as centroid distances, centroid cosine similarity, and alignment cosine similarity of representations across inter-class and intra-class sub-manifold pairs. To ensure robust estimation, we randomly sampled 800 images from the test dataset and generated 100 augmentations per image to compute the manifold properties. This process was repeated 10 times.
As shown in Fig. 4, intra-class and inter-class sub-manifolds exhibit clear differences across all metrics, indicating that manifolds of different classes are distinct in both position and orientation, reinforcing the potential for linear separability. Notably, the inter-class alignment similarity peaks strongly at $\sim 0$, suggesting that sub-manifolds of different classes are almost orthogonal to one another.
To demonstrate the emergence of category-specific manifolds, we visualize the representation space of the CIFAR-10 test dataset (without augmentation) using t-SNE  [46], a method that preserves local neighborhood structure while projecting high-dimensional data into a lower-dimensional space. As shown in Fig. 5, representations from different classes form distinct, well-separated clusters, indicating the formation of class-specific manifolds in the learned representation space. This clear clustering underscores the structured organization induced by CLAMP and demonstrates its effectiveness in producing representation that are separable in representation space.
Methods V1 $\uparrow$ V2 $\uparrow$ V4 $\uparrow$ IT $\uparrow$
  SimCLR 0.224 0.288 0.576 0.552
SwAV 0.252 0.296 0.568 0.533
Barlow Twins 0.276 0.293 0.568 0.545
BYOL 0.274 0.291 0.585 0.55
MMCR (8 views+ME) 0.270 0.311 0.577 0.554
  CLAMP (4 views) 0.258 $\pm 0.013$ 0.336 $\pm 0.017$ 0.558 $\pm 0.004$ 0.570 $\pm 0.005$
Table 3: Brain-score:We use public data from [47] for V1 and V2, and from [48] for V4 and IT to evaluate the brain-score [49, 50]. The benchmarks for all baseline methods are from [34]. In the MMCR method, ME refers to the momentum encoder.
7 Biological implications
Figure 6: The eigenspectrum of the covariance matrix for the responses of 24000 images randomly selected from the ImageNet-1K dataset follows a power law decay $\lambda\propto n^{-1.013}$. Our model is pretrained on ImageNet-1K for 100 epochs. The inset is adapted from the eigenspectrum of the stimulus-evoked activity in mouse V1 reported in [51]
Neuroscience has long served as a source of inspiration for advancements in artificial neural networks, prompting the question of whether SSL models can replicate response statistics observed in cortex. Stringer et al. [51] analyzed stimulus-evoked activity in mouse V1 (primary visual cortex) and found that the eigenspectrum of its covariance matrix follows a power-law decay, $\lambda\propto n^{-\alpha}$ with $\alpha\approx 1.04$, and further proved that $\alpha>1$ is indeed necessary to ensure differentiability of the neural code. After training CLAMP on ImageNet-1K, we measured the eigenspectrum of its feature covariance and observed a power law decay $\lambda\propto n^{-1.013}$, matching the cortical exponent and theoretical predictions [51].
We then evaluated representational alignment using the Brain-Score benchmark [49, 50], which measures how well model activations predict primate neural recordings via cross-validated linear regression. We found that CLAMP achieves the highest Brain-Score in V2 and IT compared to the selected self-supervised models. These results suggest that CLAMP ’s representation geometry not only supports downstream classification performance but also offers a compelling account of how complex visual features may self-organize in cortex without explicit supervision.
8 Discussion
We have introduced CLAMP, a novel self‑supervised learning framework that recasts contrastive learning as a neural‑manifold packing problem, guided by a physics‑inspired loss. By promoting similarity of positive pairs and repulsion of negatives through purely short-range, multi-view repulsive interactions, CLAMP achieves competitive, and in some cases state-of-the-art, performance across standard downstream benchmarks. Importantly, our analysis of training dynamics reveals a progressive structuring of the embedding space. Sub‑manifolds that are initially overlapping evolve into well-separated compact structures. We further show that class-level manifolds emerge from the aggregation of sub-manifolds, which interact purely through repulsion, a phenomenon reminiscent of phase separation in statistical physics. The mechanism underlying this emergent organization should be elucidated in future theoretical work. It is also interesting to test if hierarchical cluster emerges during the learning process in the future.
We demonstrate that CLAMP offers a compelling account of how complex visual features might self-organize in the cortex without explicit supervision, based on direct comparisons between its learned representations and experimental neural data. However, these findings should be interpreted with caution. Brain-Score’s linear mapping between deep nets output and experimental activity may overlook non-linear relationships between model and neural representation. It also does not guarantee that CLAMP implements the same computations or dynamics as cortex. Moreover, CLAMP’s use of large-batch contrastive optimization with backpropagation is unlikely to reflect biologically realistic learning mechanisms. In contrast, Hebbian-style self-supervised learning frameworks with local update rules are considered biologically plausible [52]. Developing local self-supervised learning dynamics inspired by manifold packing, without relying on backpropagation, is an important direction for future research.
Our loss function leverages the sizes and locations of sub‑manifolds within the embedding space. Incorporating orientation alignment among sub‑manifolds may enable denser packing, potentially improving downstream task performance. However, we currently omit orientation information due to its high computational cost: accurately estimating a sub‑manifold’s orientation requires on the order of $O(D)\sim 100$ points, $\sim 100$ augmentations per input, which is infeasible at the scale of CIFAR‑10 or ImageNet.
Acknowledgments
We thank Shivang Rawat, Mathias Casiulis and Satyam Anand for valuable discussions. This work was supported by National Science Foundation grant IIS-2226387, the National Institutes of Health under award number R01MH137669, the Simons Center for Computational Physical Chemistry, and NYU IT High Performance Computing resources, services, and staff expertise. We also gratefully acknowledge the use of research computing resources provided by the Empire AI Consortium, Inc., with support from the State of New York, the Simons Foundation, and the Secunda Family Foundation.
Appendix A Algorithm
Algorithm 1 CLAMP algorithm
Input: batch size $b$, number of views $m$, number of optimization steps $n$,
size scaling factor $r_{s}$
initialize feature extractor network $f_{\theta}$
initialize projection head $g_{\phi}$
initialize the distribution of random transforms $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$
for $t=1$ to $n$ do
  sample a batch from the dataset $x_{1},x_{2},...x_{b}$
  for $i=1$ to $b$ do
   # for each image, augment half of the views with $\mathcal{T}_{1}$
   $x_{i}^{1},...x_{i}^{m/2}=t^{1}(x_{i})...,t^{m/2}(x_{i}).\quad t^{1}...t^{m/2}\sim\mathcal{T}_{1}$
   #augment the other half of the views with $\mathcal{T}_{2}$
   $x_{i}^{m/2+1},...x_{i}^{m}=t^{m/2+1}(x_{i})...,t^{m}(x_{i}).\quad t^{m/2+1}...t^{m}\sim\mathcal{T}_{2}$
   $y_{i}^{1},...y_{i}^{m}=f_{\theta}(x_{i}^{1}),...f_{\theta}(x_{i}^{m})$ # extract features
   $z_{i}^{1},...z_{i}^{m}=g_{\phi}(y_{i}^{1}),...g_{\phi}(y_{i}^{m})$ # projection
  end for
  # normalization
  $c=\frac{1}{mb}\sum_{i=1}^{b}\sum_{k=1}^{m}z_{i}^{k}$
  $\tilde{z}_{i}^{k}=(z_{i}^{k}-c)/\lVert z_{i}^{k}-c\rVert_{2}$ for i=1b,k=1mi=1...b,k=1...m
  for $i=1$ to $b$ do
   $q_{i}=\frac{1}{m}\sum_{k}\tilde{z}_{i}^{k}$ # center of each sub-manifold
   $\Lambda_{i}=\frac{1}{m}\sum_{k}(\tilde{z}_{i}^{k}-q_{i})^{T}(\tilde{z}_{i}^{k}-q_{i})$ # covariance matrix
   $r_{i}=r_{s}\sqrt{\frac{1}{m}\mathrm{Tr}(\Lambda_{i})}$ # manifold size
   # Note: in practice, only the diagonal elements of $\Lambda_{i}$ are calculated for $\mathrm{Tr}(\Lambda_{i})$.
  end for
  calculate the loss function
[…]
  $\mathcal{L}=\log(\mathcal{L}_{overlap})$
  update the network $f_{\theta},g_{\phi}$ to minimize $\mathcal{L}$
end for
return $f_{\theta}$ and discard $g_{\phi}$
Appendix B Image augmentation
During self-supervised training, CLAMP employs a collection of image augmentations, drawn from the broader set introduced in [1] and [2]. We adopt the following image augmentations: randomly sized cropping, Gaussian blur, random gray-scale conversion(color dropping), random color jitter,randomly horizontal flipping.
parameter value $(\mathcal{T}_{1}=\mathcal{T}_{2})$
random sized cropping - output size 32 $\times$ 32
random sized cropping - scale [0.08,1.0]
Gaussian blur - probability 0.5
Gaussian blur - kernel size 3 $\times$ 3
color drop - probability 0.2
color jitter - probability 0.8
color jitter - brightness adjustment max intensity 0.8
color jitter - contrast adjustment max intensity 0.8
color jitter - saturation adjustment max intensity 0.8
color jitter - hue adjustment max intensity 0.2
horizontal flipping - probability 0.5
Table 4: Image augmentation parameters for CIFAR10
parameter $\mathcal{T}_{1}$ $\mathcal{T}_{2}$
randomly sized cropping - output size 224 $\times$ 224 224 $\times$ 224
randomly sized cropping - scale [0.08,1.0] [0.08,1.0]
Gaussian blur - probability 0.8 0.8
Gaussian blur - kernel size 23 $\times$ 23 23 $\times$ 23
color drop - probability 0.2 0.2
color jitter - probability 0.8 0.8
color jitter - brightness adjustment max intensity 0.8 0.8
color jitter - contrast adjustment max intensity 0.8 0.8
color jitter - saturation adjustment max intensity 0.2 0.2
color jitter - hue adjustment max intensity 0.1 0.1
horizontal flipping - probability 0.5 0.5
solarization - probability 0.0 0.2
Table 5: Image augmentation parameters for ImageNet-1K/100
As the last step of the augmentation process, we apply channel-wise normalization using dataset-specific statistics. Each image channel is standardized by subtracting the dataset mean and dividing by the standard deviation. For CIFAR-10, we use (0.4914, 0.4822, 0.4465) as the mean values and (0.247,0.243,0.261) for the standard deviation. For ImageNet-100 and ImageNet-1K, the normalization parameters are (0.485, 0.456, 0.406) for the mean and (0.229, 0.224, 0.225) for the standard deviation. During resizing, we employ bicubic interpolation to preserve image quality. We use Albumentaions library for fast image augmentation [53].
Appendix C Time complexity for the loss function
Note that only the diagonal part of the covariance matrix $\Lambda_{i}$ is required in the loss function, it is not necessary to calculate the complete covariance matrix. Therefore, each iteration, the time complexity for computing sub-manifold sizes $r_{i}$ is $O(bmD)$, the time complexity for the pairwise loss function is $O(b^{2}D)$, therefore the time complexity for each iteration is O(bDmax(m,b))O(bD\ $\text${max}(m,b)). Since the number of iteration is $N/b$ where $N$ is the size of the dataset, the complexity for calculating the loss function for one epoch is O(NDmax(m,b))O(ND\ $\text${max}(m,b)).
Appendix D Self-supervised pretraining
We employ the LARS optimizer for self‑supervised learning with its default trust coefficient (0.001). Our learning‑rate schedule begins with a 10‑step linear warm‑up, after which we apply cosine decay. To enable fully parallel training, we convert all BatchNorm layers to synchronized BatchNorm and aggregate embeddings from all GPUs when computing the loss. In addition, we exclude every bias term and all BatchNorm parameters from weight decay and use weight decay parameter = $10^{-6}$. For CIFAR-10 dataset, we removed the max pooling layer and modified the first convolution layer to be kernel size=1,stride=1,padding=2. We use $m=16$ views for CIFAR-10 dataset and $m=4,8$ for ImageNet-1K/100. $r_{s}=8.5$ is applied for CIFAR-10 and ImageNet-100/1K. Other parameters are shown in Table 6 where the base learning rate = lr $\times$ batch size/256.
dataset backbone MLP head batch size lr momentum parameter epochs
CIFAR-10 ResNet-18 [2048,256] 128 2.0 0.9 $1000$
ImageNet-100 ResNet-50 [8192,8192,512] 512 2.0 0.9 $200$
ImageNet-1K ResNet-50 [8192,8192,512] 512 1.1 0.9 $100$
Table 6: Self-supervised pretraining setups for CIFAR-10 and ImageNet-1K/100
Appendix E Linear evaluation
We discard the projection head and freeze the backbone network during linear evaluation.
CIFAR10 dataset: We use the Adam optimizer with learning rate $0.05\times$ and batch size/ 256.0 (batch size = 1024) to train the linear classifier for 100 epochs. We apply a cosine learning rate schedule.
ImageNet-100 dataset: We use SGD optimizer with Nesterov momentum and weight decay to train the linear classifier where the learning rate $0.05\times$ batch size / 256, batch size = 1024, momentum = 0.9 and weight decay = 1e-5. During training, we transform an input image by random cropping, resizing it to $224\times 224$, and flipping the image horizontally with probability 0.5. At test time we resize the image to $256\times 256$ and center-crop it to a size of $224\times 224$. We apply a cosine learning rate schedule to train the linear classifier for 100 epochs.
ImageNet-1K dataset:we use the same setup as ImageNet-100 but base learning rate $1.6\times$ batch size / 256, batch size = 1024.
Appendix F Linear evaluation on CIFAR10
Method Top1 accuracy Top5 accuracy
SimCLR 90.74 99.75
SwAV 89.17 99.68
DeepCluster V2 88.85 99.58
Barlow Twins 92.10 99.73
BYOL 92.58 99.79
CLAMP (ours) 90.21 99.60
Table 7: Linear evaluation on CIFAR10 dataset using resnet-18 as the backbone encoder. Results for SimCLR [1], SwAV  [29], DeepCluster V2 [29],Barlow twins [42] and BYOL [2] methods are taken from solo-learn benchmarking [44].
Note that we adopt the same learning rate used for ImageNet-100 and do not perform further hyperparameter tuning.
Appendix G Semi-supervised learning
We use semi-supervised learning via fine tuning on the backbone network with a linear classifier for 1% and 10% splits of the ImageNet-1K dataset. The same splits are applied as [1]. We use the SGD optimizer with momentum of 0.9 but no weight decay for cross entropy loss. We adopt the same data augmentation procedure during training and testing as in the linear evaluation protocol. We scale down the learning rate of the backbone parameters by a factor of 20. Both the backbone and classifier use the cosine-decay learning rate schedule. We use a batch size of 256 and sweep the initial learning rate over {0.1, 0.2, 0.4, 0.6}.
Appendix H Object detection
we apply the Faster R-CNN detector with a R50-C4 backbone (pretrained on ImageNet-1K with 8 views and batch size 512 for 100 epochs), fine-tuned on VOC2007+2012 training dataset and tested on VOC2007 test dataset. Our implementation follows the procedure of [3], except that we set the base learning rate to 0.07.
Appendix I Power law exponent for the eigen-spectrum
In analysis of experimental data in [51], the exponent is estimated using log-log scale liner fit on eigenvalues ranked from 11 to 500. Here, we use eigenvalues ranked from 15 to 1400 to fit the exponent.
Appendix J Brain-Score
We use the datasets with identifiers, FreemanZiemba2013public.V1-pls,
FreemanZiemba2013public.V2-pls, MajajHong2015public.V4-pls, MajajHong2015public.IT-pls for the neural activity recorded from V1,V2,V4 and IT respectively.
Appendix K Qualitative understanding of sub-manifolds in the embedding space
K.1 Manifold radius for singular $\Lambda_{i}$
As $\Lambda_{i}$ is positive semi-definite, it can be decomposed as
$\Lambda_{i}=\sum_{j}^{K}\lambda_{i}^{(j)}v_{i}^{(j)}v_{i}^{(j)^{T}}$
with $\lambda_{i}^{(k)}>0$ where $K=rank(\Lambda_{i})$ and $v_{i}^{(j)}$ are orthonormal vectors. We then enclose the sub-manifold as the ellipsoid whose equation reads,
$\frac{[(\tilde{z}-Z_{i})^{T}v_{i}^{(1)}]^{{}^{2}}}{r_{s}^{2}\lambda_{i}^{(1)}}+\frac{[(\tilde{z}-Z_{i})^{T}v_{i}^{(2)}]^{{}^{2}}}{r_{s}^{2}\lambda_{i}^{(2)}}+...+\frac{[(\tilde{z}-Z_{i})^{T}v_{i}^{(K)}]^{{}^{2}}}{r_{s}^{2}\lambda_{i}^{(K)}}=1.$
This equation represents the K-dimensional ellipsoid centered at $Z_{i}$ embedded in the D-dimensional space with semi-axial lengths $r_{s}\sqrt{\lambda_{i}^{(1)}}...r_{s}\sqrt{\lambda_{i}^{(K)}}$. We compute the sub-manifold radius as the square root of average semi-axial length square,
$r_{i}=\sqrt{\frac{r_{s}^{2}\lambda_{i}^{(1)}+...+r_{s}^{2}\lambda_{i}^{(K)}}{K}}.$
Even if $\Lambda_{i}$ is not a full-rank matrix, $\mathrm{Tr}(\Lambda_{i})=\lambda_{i}^{(1)}+...+\lambda_{i}^{(K)}$ is well defined. We get
$r_{i}=r_{s}\sqrt{\frac{\mathrm{Tr}(\Lambda_{i})}{K}}$
$\Lambda_{i}$ is the covariance matrix of $m$ embedding vectors embedded in D-dimensional space, Therefore $rank(\Lambda_{i})\leq min(m,D)=m$ (in practice, the number of views is much smaller than the dimensionality , $m\ll D$). Here, we use $m$, the upper bond, to estimate $rank(\Lambda_{i})$.
K.2 The relation between estimated sub-manifold radius and the estimated sub-manifold volume
We enclose the sub-manifold $i$ by $K$ ($K=rank(\Lambda_{i})$) dimensional ellipsoid embedded in the D-dimensional space with semi-axial lengths $r_{s}\sqrt{\lambda_{i}^{(1)}}...r_{s}\sqrt{\lambda_{i}^{(K)}}$. Its volume reads
$V=\frac{\pi^{K/2}}{\Gamma(\frac{K}{2}+1)}\prod_{j}^{K}r_{s}\sqrt{\lambda_{i}^{(j)}}.$
From the inequality of arithmetic and geometric means, we have
$V\leq\frac{\pi^{K/2}}{\Gamma(\frac{K}{2}+1)}\left(\frac{1}{K}\sum_{i}{r_{s}\sqrt{\lambda_{i}^{(j)}}}\right)^{K}$
By Cauchy–Schwarz inequality
$V=\frac{\pi^{K/2}}{\Gamma(\frac{K}{2}+1)}\left(\frac{1}{K}\sum_{i}{r_{s}\sqrt{\lambda_{i}^{(j)}}}\right)^{K}\leq\left(r_{s}\sqrt{\frac{\mathrm{Tr}(\Lambda_{i})}{K}}\right)^{K}\frac{\pi^{K/2}}{\Gamma(\frac{K}{2}+1)}=\frac{\pi^{K/2}}{\Gamma(\frac{K}{2}+1)}r_{i}^{K}$
Therefore $r_{i}$ is related to the upper-bound of the estimated sub-manifold volume.
Appendix L Hyperparameters
We conducted an ablation study to investigate the role of different hyperparameters. Due to limited computational resources, ablation study on parameter $r_{s}$ and number of views $m$, $30\%$ of the training data from ImageNet-1K were selected and pretrained for 100 epochs (which is consistent with our baseline where the whole training set are pretrained for 100 epochs). For the ablation on batch size, we employed the entire ImageNet-1K dataset. We used the frozen-backbone linear evaluation accuracy on the entire test dataset as the criterion for ablation.
L.1 Scale parameter $r_{s}$
As shown in Fig. 7, accuracy increases with $r_{s}$ until it reaches a plateau at approximately $r_{s}\approx 6.0$. This result further supports our qualitative finding that setting $r_{s}>3$ is necessary to compensate for the underestimation of manifold size. A higher $r_{s}$ increases computational cost because more sub-manifold pairs are involved in the loss function and the corresponding optimization process.
Figure 7: Linear evaluation accuracy as the function of the size scale factor $r_{s}$ on $30\%$ ImageNet-1K dataset. we use 4 views.
L.2 Number of augmentations $m$
We observed that increasing the number of data augmentations leads to higher linear evaluation accuracy. This indicates that greater augmentation diversity helps the model learn better representations.
2 views 4 views 8 views
top 1 accuracy 55.17 58.29 58.95
Table 8: Linear evaluation using different number of views for ablation study on $30\%$ ImageNet-1K dataset
L.3 Batch sizes
In contrast to SimCLR, where large batch sizes are critical for achieving high accuracy, we find that the linear evaluation accuracy of CLAMP depends only weakly on batch size.
batch size 256 512 1024
top 1 accuracy 69.02 69.43 69.11
Table 9: Linear evaluation using different batch sizes for 4 views
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims presented in the abstract and introduction are substantiated by our methodology, evaluation results, and supporting analyses, including figures and tables.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss our limitation in the discussion section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Please see the section method and the Appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We show the details of the model, parameters, datasets in section implementation details and appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We include our code in the supplementary material and all the datasets we use are publicly available.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We have presented the benchmark and implementation details of our algorithms in in section implementation details and appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report the error (two-sigma for three independent experiments with different random seeds) in the table. The error in Brain-Score is provide by the library.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We have these details in section implementation details.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our research conforms in every points with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss the broader impacts in our abstract and introduction.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification:This paper has no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification:All the assets are properly credited and respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: No new assets are created in this paper.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve any crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve any crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.